From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 15 2003 - 23:33:32 EST
------------------------------------------------------------ Date: Sat Feb 15 23:33:31 2003, Shift: 16:00-24:00 Supervisor: ZY, with: Mads Log book#14 pages 19 - 19 Shift Summary: ------------------------------------------------------------ 1. RHIC magnet problems, no beam, after the daq was fixed around six. 2. There was beam at the begining of the shift but data could not be taken because of daq problem. (see 3) No data taken in shift. Pedestal runs and DVM runs. 3. At begining of shift there were problems with the daq, fixed by JH Lee, I refer to the Elog. Brahms detectors seem to be functioning normally. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sat Feb 15 2003 - 23:36:16 EST